OpenStudy (anonymous):

Juliana has created the function f(x)=3x+2/4 to represent the cost of texting on her current plan, where x represents the number of texts. Juliana discovers that using the inverse function to solve for x=24, she can predict how many texts she can use for $24. Explain to Juliana how to accomplish this, using complete sentences

OpenStudy (anonymous):

@Gerardo_cast23 @Falco276 @Hero

hero (hero):

Is it \[f(x) = \frac{3x + 2}{4}\] or \[f(x) = 3x + \frac{2}{4}\]

OpenStudy (anonymous):

the first one!

hero (hero):

What's interesting is, you actually have to find the inverse of the function.

hero (hero):

I suppose that's what you're having trouble with, right?

OpenStudy (anonymous):

YES exactly

hero (hero):

To find the inverse of the function, you're supposed to change \(f(x)\) to \(y\), swap \(x\) and \(y\), then isolate y afterwards.

hero (hero):

So you have \[f(x) = \frac{3x + 2}{4}\] Changing \(f(x)\) to \(y\) you get: \[y = \frac{3x + 2}{4}\] Swapping \(x\) and \(y\) you get: \[x = \frac{3y + 2}{4}\] Can you isolate \(y\) from here?

OpenStudy (anonymous):

thats where I have the problemm

hero (hero):

To finish, 1. Multiply 4 on both sides 2. Subtract 2 on both sides 3. Divide both sides by 3

OpenStudy (anonymous):

i dont get where u multiply the 4 on the fraction side

hero (hero):

\[4 \dot\ x = \frac{3y + 2}{4} \dot\ 4\]\[4 \dot\ x = \frac{3y + 2}{\cancel{4}} \dot\ \cancel{4}\]\[4x = 3y + 2\]

OpenStudy (anonymous):

y = 4x-2 / 3

hero (hero):

Lastly you have to replace \(y\) with \(f^{-1}(x)\)

OpenStudy (anonymous):

f -1 (x) = y= 4x-2 /3

hero (hero):

Now, that you have inversed the function, the domain and range are the inverse of what they were before.

OpenStudy (anonymous):

Thank you! Can u help with another pleasE?

OpenStudy (anonymous):

oh but now my answer..do i sub 24?

hero (hero):

I was just about to ask you about that

hero (hero):

You have to find \(f^{-1}(24)\)

OpenStudy (anonymous):

So 4(24) -2 / 3?

OpenStudy (anonymous):

@Hero

hero (hero):

Yes, what did you get?

OpenStudy (anonymous):

31.3

hero (hero):

Well, we should say approximately 31 texts for $24

hero (hero):

Since it is not possible to send .3 texts.

OpenStudy (anonymous):

lol yeah ok so 31 texts she can send

hero (hero):

Correct.

hero (hero):

To check, input f(31) into the original function to see if you get 24.

OpenStudy (anonymous):

Ok . please help me understand this one

OpenStudy (anonymous):

Marrie is logging the number of radioactive isotope she observes every minute. She has determined her function to be f(x) =-4x +120. Based on the situatuion above describe and neccesary restrictions to the domain and range

hero (hero):

I've answered this kind of question already. You should scroll through my questions answered to find it.

OpenStudy (anonymous):

okkay !

OpenStudy (anonymous):

u have way too many lol

OpenStudy (anonymous):

Thank you @broskishelleh

OpenStudy (anonymous):

@Hero You're a saint! Thanks for clearing that up!